OpenStudy (anonymous):

Series : converges or not : ((cos n)^2)/(n^2 + 1) ? and how you do it ?

OpenStudy (anonymous):

Use the squeeze theorem. Since:\[0\le \left(\cos n \right)^2\le 1\]Dividing everything by n^2+1 yields:\[0\le \frac{(\cos n )^2}{n^2+1}\le \frac{1}{n^2+1}\]What is the limit of the rightmost term as n goes to infinity? What is the limit of the leftmost term? What can you conclude about the middle term?

OpenStudy (anonymous):

but the left one is not -1 / (n^2 +1) ?

OpenStudy (anonymous):

Generally that is the way to go, but the square around the cos n term makes it positive again.

OpenStudy (anonymous):

ooo alright thx alot !

OpenStudy (anonymous):

the limit is 0 ?

OpenStudy (anonymous):

Yes thats correct :)

OpenStudy (anonymous):

but converges or not because general term say if = 0, we cant conclude anything

OpenStudy (anonymous):

Ah, I didnt realize it was a question about a series, and not a sequence, my apologies.

OpenStudy (anonymous):

So you are correct, that test doesn't conclude anything about the series. So we should use a direct comparison.

OpenStudy (anonymous):

Note that:\[\frac{(\cos n )^2}{n^2+1}\le \frac{1}{n^2+1}\le \frac{1}{n^2}\]

OpenStudy (anonymous):

P series ? \[n > 1\] so converges ?

OpenStudy (anonymous):

Thats correct, so since \[\sum_{n=1}^{\infty}\frac{1}{n^2}\]converges, and:\[\frac{(\cos n)^2}{n^2+1}\le \frac{1}{n^2}\], what can we conclude about:\[\sum_{n=1}^{\infty}\frac{(\cos n)^2}{n^2+1}\]?

OpenStudy (anonymous):

it converges !